OpenStudy (anonymous):

Find the inverse of each function. Is the inverse a function? 1) f(x)=(x+2)^2-4 2) f(x)=4x^3-1 3) f(x)=square root of x+4

OpenStudy (anonymous):

Okay so for the first one, you want to set: y=(x+2)^2-4 then change the y to x, and change the x-terms to y. x=(y+2)^2-4 Then solve for y. You'll have y = something. Then just replace y with f(x).

OpenStudy (anonymous):

sqrt(x+4) = y+2 y=sqrt(x+4)-2 f(x) = y=sqrt(x+4)-2

OpenStudy (anonymous):

how do i get y by itself

OpenStudy (anonymous):

I meant to say: f(x) = sqrt(x+4)-2

OpenStudy (anonymous):

"how do i get y by itself" You have to square root both sides.

OpenStudy (anonymous):

x=(y+2)^2-4 x+4 = (y+2)^2 sqrt(x+4) = sqrt((y+2)^2) sqrt (x+4) = y+2 y = sqrt(x+4) -2

OpenStudy (anonymous):

the -2 is outside the square root?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

so that means its not a function right, since it has a square root?

OpenStudy (anonymous):

uh... not quite. It's just not continuous, where the stuff inside the square root equals a negative number. It's a "discontinuous" function.

OpenStudy (anonymous):

So anything where x is less than -4 (like... x<-4, -5,... infinity), f(x) is undefined.

OpenStudy (anonymous):

wait so the inverse is a function or not?

OpenStudy (anonymous):

I'm sorry, I should have referred to the inverse as f^-1(x). So anything where x is less than -4 (like... x<-4, -5,... infinity), f^-1(x) is undefined. f^-1(x) =sqrt(x+4)-2 This is the inverse function.

OpenStudy (anonymous):

so it is a function?

OpenStudy (anonymous):

Yes, it is a function. Inverse functions are functions... Just like functions are functions :-)

OpenStudy (anonymous):

Here's a link that should help you more http://tutorial.math.lamar.edu/Classes/CalcI/InverseFunctions.aspx

OpenStudy (anonymous):

ok for number 2 you cube root?

OpenStudy (anonymous):

Yup! Show me your process and I'll check it.

OpenStudy (anonymous):

i got cube root of x=1/4=y

OpenStudy (anonymous):

x+1

OpenStudy (anonymous):

first i added 1, cube rooted, and divided by 4

OpenStudy (anonymous):

x=4y^3-1 x+1 = 4y^3 (x+1)/4 = y^3 (x/4+1/4) = y^3 cubedroot(x/4+1/4) = y

OpenStudy (anonymous):

is cube root of x+1/4 ok?

OpenStudy (anonymous):

They both need to be over 4. B/c when you divide on one side, then you divide everything by that number on the other side.

OpenStudy (anonymous):

\[\sqrt[3]{x+1}/4\]

OpenStudy (anonymous):

the 4 needs to be inside of the cubed root. cubed root of (x/4 +1/4)

OpenStudy (anonymous):

can you do it using the equation thing please

OpenStudy (anonymous):

\[\sqrt[3]{(x/4)+(1/4)}\]

OpenStudy (anonymous):

so its not the same thing as everything over 4?

OpenStudy (anonymous):

(x/4) + (1/4) = (x+1)/4

OpenStudy (anonymous):

So what I did in the equation thing would be the same thing as saying: \[\sqrt[3]{((X+1)/4)}\]

OpenStudy (anonymous):

that wouldn't be a function right

OpenStudy (anonymous):

It would. Just not any place where stuff inside the cubed root would equal a negative number.

OpenStudy (anonymous):

how do you determine if its a function or no?

OpenStudy (anonymous):

Where x=-1 (-1+1)/4 = 0/4 (this IS defined, but anything less than -1 is undefined) Where x=-1.5 (-1.5+1)/4 = -.5/4 The cubed root of -.5/4 is undefined.

OpenStudy (anonymous):

undefined would mean not a function?

OpenStudy (anonymous):

I'll research that, I'm sorry for not answering the second part of the question very well. One sec.

OpenStudy (anonymous):

for the last one i got x^2-4 = y

OpenStudy (anonymous):

As far as I know, if the inverse of a function only has one output per each input, then it is a function. Try going onto google and graph x^(1/2) and x^(1/3). Those are similar functions to what you have in your answers for a. and b. so with those, you'll see that are functions. A problem you usually run in to is when you see something like y=sqrt(x^2 +2). That is a circle, so it is not a function. You don't have anything to worry about at least with the first and second ones.

OpenStudy (anonymous):

Paul Dawkins give a really good set of instructions on how to find an inverse and then how to PROVE that what you found is truly the inverse of your function: http://tutorial.math.lamar.edu/Classes/Alg/InverseFunctions.aspx

OpenStudy (anonymous):

Let's see... x^2 = y+4 y = x^2 -4 Yep, you are right on that last one. And it is a function.

OpenStudy (anonymous):

the first 2 don't look like functions from what i remember so im not sure

OpenStudy (anonymous):

sometimes google graphs things wrong. I will go bust out my calculator

OpenStudy (anonymous):

Try graphing the first and second ones. got a graphing calculator?

OpenStudy (anonymous):

This is an example of something that is NOT a function.

OpenStudy (anonymous):

so im guessing they're not funtions

OpenStudy (anonymous):

I keep getting conflicing answers. Must be that anything with a radical is not a function. Sorry for steering you wrong :-(

OpenStudy (anonymous):

it's okay, how would something with a radical's graph look anyways?

OpenStudy (anonymous):

I have 3 more questions if you dont mind helping me

OpenStudy (anonymous):

Well I guess you would graph the f(x)= +sqrt(x+4)-2 and f(x)= -sqrt(x+4)-2 http://www.ehow.com/how_4515289_square-root-functions-fx-x.html

OpenStudy (anonymous):

I'll do the rest tomorrow, thanks for helping!

OpenStudy (anonymous):

k :-) I'll look into the one about the cubed root.. I'm a little unsure on that one

OpenStudy (anonymous):

Looks like with anything that has a radical which is an odd number... like the cubed root, the 5th root, 7th root, whatever...that those are functions. Here's what I'm thinking. If you have the sqrt(4), then what are the possible answers? Well (-2)(-2) =4 AND (2)(2) =4. So square roots or any root that is an even number can't possibly be a function. Okay well if you have the CUBED root(8), then what is/are the possible solution(s)? (2)(2)(2) = 8 BUT! You can't have: (-2)(-2)(-2) does not equal 8. Therefore, cubed roots (or any root that is an odd number) only have one solution and is a function. Square roots(or any root that is an even number) always has two solutions and is NOT a function. This was a good question, it helped refresh my memory on what is/isn't a function :-)

OpenStudy (anonymous):

Oh and here's a video... he clarifies what I just said really well: http://www.youtube.com/watch?v=oJLLQ2WeT5M